what is the solution in this problem?The ratio between the base and the altitude is 2:3.If the area of the triangle is 108 sq. cm.,what is the base?
Answer by mananth(16946)

let the common multiple be x
height = 2x
base = 3x
Area = 1/2 * base * height
108= 1/2 * 2x * 3x
3x^2=108
x^2=36
x= +6
height = 2x= 12
base = 3x= 18